Suppose 207 subjects are treated with a drug that is used to treat pain and 54 of them developed nausea. Use a 0.05 significance level to test the claim that more than 20​% of users develop nausea.
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p =0.20

Ha : p >0.20

= x / n = 207 /54

P0 = 0.20

1 - P0 = 1 - 0.20 =0.80

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.2609 - 0.20/ [(0.20*0.80) / 207]

= 2.19

P(z >2.19 ) = 1 - P(z <2.19 ) = 1 - 0.9857 =0.0143

P-value = 0.0143

= 0.05

0.0143 < 0.05

it is concluded that the null hypothesis is rejected.

Conclusion

It is concluded that the null hypothesis Ho is rejected. Therefore, there is enough evidence to claim that the population proportion p is greater than p0, at the α=0.05 significance level.